How much of a solution that is 30% alcohol must be added to 55 liters of a solution that is 60% alcohol to obtain a solution that is 40% alcohol?
Answer by TimothyLamb(4379)

30w + 60(55) = 40(w + 55)
30w + 3300 = 40w + 2200
10w = 1100
---
w = 110 L
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php